# Large Language Models as Realistic Microservice Trace Generators

### v1

@@ 4.5 Adapting Models to Solve Downstream Tasks
The second experiment’s dataset comprises 1K samples, each representing a pair of parent and child layers with a missing connecting edge tagged as [MISSING]. After training, we test both models on 5K test cases to generate the correct edge, ensuring the finish time matched or exceeded the start time. The right part of Figure 7 shows that while the original Llama-2 model scored only 24% accuracy and Llama-3.1 405B reached 34%, our model maintained a high accuracy of 66%, underscoring its robustness in more complex tasks.
@@
5 Other Related Work
@@
Making Language Models Follow Instructions.  Recent advancements have focused on enhancing LLMs’ ability to follow instructions through prompting (Li & Liang, 2021; Shin et al., 2020; Wei et al., 2022) and instruction tuning (Ouyang et al., 2022; Wei et al., 2021; Chung et al., 2022). These two sets of methods are relevant to our setting since they augment powerful pre-trained LLMs to improve their performance on new tasks. Our approach seeks to refine output expressiveness within set prompts, aiming for greater fidelity in synthetic data production.
@@
Multi-step Reasoning with LLMs.  Iterating with LLMs over multiple steps is an effective strategy to solve complex problems. For instance, Tree-of-thoughts (Yao et al., 2024) solves problems by decomposing into smaller thoughts and exploring diverse reasoning paths over different thoughts. Multi-step reasoning is also useful to handle long-context scenarios by summarizing iteratively (Wang et al., 2023) and diving into subproblems (Lee & Kim, 2023). In contrast to the above approaches, our approach learns to generate traces with specific structures and instructions for subsequent layers.
6 Conclusion
This paper presents a training method for pre-trained LLMs tailored for generating microservice trace graphs through a recursive call graph generation scheme complemented by instruction tuning with intermediate instructions. Our evaluation results demonstrate that our approach outperforms the baselines in generating accurate and valid call graphs and shows improved distributional similarity to real-world traces. In addition, we show that synthetic traces can effectively serve as a substitute for real data in training microservice management tasks, such as critical component detection and anomaly detection. They further highlight the effectiveness of instruction tuning in refining the generation of call graphs according to user-specified features and reveal the potential for using our model in various downstream tasks, such as prediction and data infilling, by further training the model. While this paper focuses primarily on microservice call graphs, our approach holds promise for broader applicability to other computer system traces with similar structural characteristics.
Ethics Statement.  There are no ethical concerns raised by our work as the data used in this study is public and sensitive information has been anonymized.
Reproducibility Statement.  We include the model and datasets used in experiments along with the steps to preprocess datasets. We describe training hyperparameters and data preprocessing steps in Appendix A. We include scripts to reproduce our experiments in the supplementary material.
Appendix A Training Details
A.1 Training Setup
@@
We train all models with 4$\times$ A100 80GB GPUs in our cluster with the hyperparameters described in Table 1. We apply LoRA ((Hu et al., 2022)) adapters to query and key projection matrices of attention layers with $rank=8$, $alpha=16$, and $dropout=0.1$. For the downstream task training in § 4.5, we freeze the backbone model and only train the last classification layer for the prediction task. For the infilling downstream task, we use LoRA adapters with the same configuration as mentioned earlier.
A.2 Training Data Preprocessing
From the Alibaba microservice v2022 traces (Luo et al., 2022), we use the first-hour call graph traces as our training data, which consist of 6434 unique microservices collected from more than 10 clusters. The traces are collections of API calls, where each API call includes communication information between the two microservices. Note that the dataset anonymizes the service and microservice names. Service ID is a nine-digit number starting with the prefix "S_" instead of using a real service name (e.g., social network), and microservice is a five-digit number starting with the prefix "MS_". We construct call graphs using the trace ID field (i.e., API calls with the same trace ID belong to one call graph). When constructing call graphs, we remove calls with missing information (e.g., destination microservice IDs are unknown) and remove call graphs that are not connected (e.g., missing edges). To remove redundancy, we filter out call graphs that have the same structure and fields (e.g., service ID, latency) for all API calls. The distributions of training data after removing redundancy are shown in Figure 8.
A.3 Training Data Examples
From the call graph traces, we create text-based representations of call graphs as described in Section 3.1.1. First of all, Figure 9 is a training data example of converting a call graph into a tabular data format, which is the baseline in § 4.1. At the beginning, we include high-level information about the call graph including service ID, the number of edges, and depth of the call graph. Each line inside the <edges> block corresponds to a call in a call graph. 6 fields exist for each call including the edge ID, source/destination microservices, communication type, and communication start/finish time.
Figure 11 is an example of instruction-tuning data. The instruction starts with a system prompt followed by conditions as in Figure 10. We further explain the condition in natural language formats along with user-requested features as studied in § 4.4. In the output section, we include Chain-of-Thought scratchpads at the end of <edges> block and at the beginning of <subgraph> blocks, which elaborate on the number of edges to generate and constraints of subgraph conditions. For instance, the scratchpad includes descriptions regarding the depth requirement to let LLMs understand better that the depth field should be decreased by 1 from the current layer’s depth.
@@ A.3 Training Data Examples
Figure 8: Training data distribution after preprocessing steps.
Table 1: Training setup and hyperparameters.
@@
Pre-Training & Instruction Tuning Optimizer AdamW (Loshchilov & Hutter, 2017)
@@
Appendix B Constraints in Call Graph Layers
@@
Edges.  For each edge, we check the following conditions. First of all, each edge should include the 5 fields: edge ID, destination, communication type, and communication start/finish time. Secondly, we check whether the right number of edges are generated as described in the condition. Third, the communication start time should be equal to or greater than the communication start time described in the condition, and should not be greater than the communication finish time of the edge. Lastly, the communication finish time should be equal to or less than the latency field in the condition.
@@
Then, we check the validity of each field in the next layer conditions. First of all, the edge ID inside the <subgraph> block should be found in the edges generated in the current layer. For the depth, the remaining depth field should be less than the remaining depth of the instruction. Additionally, at least one of the resulting subgraphs must have a depth that is reduced by one compared to the original graph. For the start node and caller fields, they should be copied from the destination from the parent edge and the start node from the instructions, respectively. Lastly, we check the latency and communication start time by comparing the values to those of the parent edge. The latency of a child layer should not be greater than the communication finish time of the parent edge. Also, the communication start time of a child layer should not be less than the communication start time of the parent edge.
@@
Appendix C Limitations
In this section, we discuss a few limitations of our work and potential approaches to overcome the limitations. The recursive method improves call graph generation accuracy compared to generating the entire trace at once, but a key drawback is that previously generated edges are discarded, as only the conditioning information from the prior layer is passed to the next layer generation steps. Although dropping previously generated results has little impact on the output in microservice call graph generation, where direct neighbors exert the most influence (Zhang et al., 2024), we believe incorporating past information, such as previous layers or a time series of call graph traces, could be beneficial.
Furthermore, our method uses manually constructed instruction templates, which may lead to suboptimal generation quality, as we are not using the full potential of language models pre-trained with trillions of tokens (Touvron et al., 2023). Following the methods of Liu et al. (2024); Gunasekar et al. (2023); Li et al. (2024), we believe that diversifying instructions using LLM-generated output is a potential method to improve the ability of LLMs to follow user intentions.
Lastly, while protecting sensitive data during synthetic trace generation is an important research challenge, this paper does not address privacy concerns. Our framework assumes that the data used to train a model such as ours is curated at the source to remove sensitive attributes/values. As part of our future work, we plan to investigate whether our model exposes sensitive information and implement privacy-preserving techniques (e.g., differential privacy) during fine-tuning LLMs (Yu et al., 2022).
@@ D.1 Structured Reasoning Results in Detail
This section provides a more detailed analysis of the results from § 4.1, accuracy to generate call graphs adhering to all structural constraints while matching the specified attributes in prompts (i.e., num_edges and depth). Figure 15 offers a closer look at 3(a) and 3(b), where each grid point $(X,Y)$ represents accuracy for prompts with $X$ edges and a maximum depth of $Y$. 14(a), 14(b), and 14(c) correspond to the same settings as (baseline), (recursive), and (recursive + instruction) from § 4.1, respectively. The results in Figure 15 show that the recursive generation and instruction tuning improves accuracy across most combinations of (# Edges, Depth). However, some configurations in 14(b) and 14(c) exhibit lower accuracy, likely due to the distribution of training data in terms of edge count and depth.
D.2 Structured Reasoning Results varying Model Sizes
Figure 12 presents the microservice call graph generation accuracy across different model sizes. Overall, models with a larger number of parameters demonstrate higher accuracy, with this trend being particularly evident in 12(c). Notably, models with more parameters perform better as the depth of prompts increases. For instance, the 13B model achieves a 20 percentage point improvement over the 7B model for inputs with a depth greater than 4 as shown in 12(b).
@@ D.2 Structured Reasoning Results varying Model Sizes
D.3 More Experiments on Similarity Between Real and Synthetic Traces
@@
D.4 More Experiments on Using Synthetic Traces in ML Use Cases
(a) Baseline accuracy heatmap.
(b) Accuracy heatmap with recursive generation.
Figure 14: Accuracy heatmap.

### v2

@@ 4.5 Adapting Models for Trace-Related Downstream Tasks
The second experiment’s dataset comprises 1K samples, each representing a pair of parent and child layers with a missing connecting edge tagged as [MISSING]. After training, we evaluate both models on 5K test cases to generate the correct edge, ensuring the finish time matched or exceeded the start time. The right part of Figure 7 shows that while the original Llama-2 model scored only 24% accuracy and Llama-3.1 405B reached 34%, our model maintained a high accuracy of 66%, underscoring its robustness in more complex tasks.
@@
5 Conclusion
This paper introduces a training method for adapting pre-trained LLMs to generate microservice trace graphs using recursive call graph generation and instruction tuning. Our approach outperforms baselines in producing accurate, valid call graphs with improved distributional similarity to real traces. We demonstrate that synthetic traces can effectively replace real data for training microservice management tasks, such as critical component extraction and anomaly detection. Additionally, instruction tuning enhances graph generation based on user-specified features, enabling applications in prediction and data infilling. While we focus on microservice call graphs, our method broadly applies to other system traces with similar structures.
6 Limitations
This section discusses a few limitations of our work and potential approaches to overcome them.
The recursive method improves the accuracy of call graph generation compared to generating the entire trace at once, but a key drawback is that previously generated edges are discarded, as only the conditioning information from the prior layer is passed to the next layer generation steps. Although dropping previously generated edges has little impact on the output in microservice call graph generation, where direct neighbors exert the most influence (Zhang et al., 2024), incorporating past information, such as prior layers or a time series of call graph traces, could enhance the capture of longer-range dependencies and temporal patterns. However, efficiently compressing historical trace information while preserving critical details remains an open challenge. In future work, we will consider this direction to compress long-range traces and generate synthetic traces conditioned on the compressed traces.
Furthermore, our method uses manually constructed instruction templates, which may lead to suboptimal generation quality, as we are not using the full potential of language models pre-trained with trillions of tokens (Touvron et al., 2023). Following the methods of prior work Liu et al. (2024); Gunasekar et al. (2023); Li et al. (2024), we believe that diversifying instructions using LLM-generated output is a potential method to improve the ability of LLMs to follow user intentions. However, naively guiding LLMs to generate instructions for trace generation may result in instructions that lack useful characteristics for downstream tasks. In future work, we plan to integrate domain-specific knowledge of traces to improve the usefulness and diversity of instructions generated by LLMs.
Lastly, we focused on generating microservice call graphs in this paper, but other system traces, such as operating system (OS) call graphs, share a similar hierarchical structure. The primary differences in OS call graphs lie in their greater depth and the increased diversity of node and edge types. Evaluating whether our approach can be effectively applied to such traces remains an open question.
Ethics Statement.  There are no ethical concerns raised by our work as the data used in this study is public with sensitive information redacted.
Appendix A Other Related Work
@@
Making Language Models Follow Instructions.  Recent advancements have focused on enhancing LLMs’ ability to follow instructions through prompting (Li and Liang, 2021; Shin et al., 2020; Wei et al., 2022) and instruction tuning (Ouyang et al., 2022; Wei et al., 2021; Chung et al., 2022). These two sets of methods are relevant to our setting since they augment powerful pre-trained LLMs to improve their performance on new tasks. Our approach seeks to refine output expressiveness within set prompts, aiming for greater fidelity in synthetic data production.
@@
Multi-step Reasoning with LLMs.  Iterating with LLMs over multiple steps is an effective strategy to solve complex problems. For instance, Tree-of-thoughts (Yao et al., 2024) solves problems by decomposing into smaller thoughts and exploring diverse reasoning paths over different thoughts. Multi-step reasoning is also useful to handle long-context scenarios by summarizing iteratively (Wang et al., 2023) and diving into subproblems (Lee and Kim, 2023). In contrast to the above approaches, our approach learns to generate traces with specific structures and instructions for subsequent layers.
Appendix B Training Details
B.1 Training Setup
@@
We train all models with 4$\times$ A100 80GB GPUs in our cluster with the hyperparameters described in Table 1. We apply LoRA ((Hu et al., 2022)) adapters to query and key projection matrices of attention layers with $rank=8$, $alpha=16$, and $dropout=0.1$. For the downstream task training in § 4.5, we freeze the backbone model and only train the last classification layer for the prediction task. For the infilling downstream task, we use LoRA adapters with the same configuration as mentioned earlier. During inference, we use a temperature of 0.8 and top-K of 50 for trace generation, unless otherwise specified. We use models (Llama 2 and 3.1) under appropriate community license.
@@
Figure 8: Training data distribution after preprocessing.
@@
Pre-Training & Instruction Tuning Optimizer AdamW (Loshchilov and Hutter, 2017)
@@
Table 1: Training setup and hyperparameters.
B.2 Training Data Preprocessing
From the Alibaba microservice v2022 traces (Luo et al., 2022), we use the first-hour call graph traces as our training data, which consist of 6434 unique microservices collected from more than 10 clusters. We reserve 10% of these samples for validation. The traces are collections of API calls, where each API call includes communication information between the two microservices. Service ID is a nine-digit number starting with the prefix "S_" instead of using a real service name (e.g., social network), and microservice is a five-digit number starting with the prefix "MS_". We construct call graphs using the trace ID field (i.e., API calls with the same trace ID belong to one call graph). When constructing call graphs, we remove calls with missing information (e.g., destination microservice IDs are unknown) and remove call graphs that are not connected (e.g., missing edges). To remove redundancy, we filter out call graphs that have the same structure and fields (e.g., service ID, latency) for all API calls. The distributions of training data after removing redundancy are shown in Figure 8.
Note that the dataset anonymizes service and microservice names, ensuring our model does not disclose sensitive information. However, if the training dataset contains sensitive data, models trained with our method may still risk exposing privacy-related information.
The instruction-tuning datasets were created by randomly selecting 5% of the training graphs and reformatting them for instruction tuning. The instruction-tuning lasted four epochs.
B.3 Training Data Examples
From the call graph traces, we create text-based representations of call graphs as described in Section 3.1.1. First of all, Figure 9 is a training data example of converting a call graph into a tabular data format, which is the baseline in § 4.1. At the beginning, we provide high-level information about the call graph, including the service ID, number of edges, and graph depth. Each line inside the <edges> block corresponds to a call in a call graph. 6 fields exist for each call including the edge ID, source/destination microservices, communication type, and communication start/finish time.
Figure 11 is an example of instruction-tuning data. The instruction starts with a system prompt followed by conditions as in Figure 10. We further explain the condition in natural language along with user-requested features, as studied in § 4.4. In the output section, we include Chain-of-Thought scratchpads at the end of <edges> block and at the beginning of <subgraph> blocks, which elaborate on the number of edges to generate and constraints of subgraph conditions. For example, the scratchpad includes descriptions of the depth requirement to help LLMs understand that the depth field should be reduced by 1 from the current layer’s depth.
(a) Baseline accuracy heatmap.
(b) Accuracy heatmap with recursive generation.
Figure 12: Accuracy heatmap.
Appendix C Constraints in Call Graph Layers
@@
Edges.  For each edge, we check the following conditions. First of all, each edge should include the 5 fields: edge ID, destination, communication type, and communication start/finish time. Secondly, we check whether the right number of edges are generated as described in the condition. Third, the communication start time should be equal to or greater than the communication start time described in the condition, and should not be greater than the communication finish time of the edge. Lastly, the communication finish time must not exceed the latency specified in the condition.
@@
Then, we check the validity of each field in the next layer conditions. First of all, the edge ID inside the <subgraph> block should be found in the edges generated in the current layer. For the depth, the remaining depth field should be less than the remaining depth of the instruction. Additionally, at least one of the resulting subgraphs must have a depth that is reduced by one compared to the original graph. For the start node and caller fields, they should be copied from the destination from the parent edge and the start node from the instructions, respectively. Lastly, we check the latency and communication start time by comparing the values to those of the parent edge. The latency of a child layer should not be greater than the communication finish time of the parent edge. Also, the communication start time of a child layer must not be earlier than that of the parent edge.
@@ D.1 Structured Reasoning Results in Detail
This section provides a more detailed analysis of the results from § 4.1, accuracy to generate call graphs adhering to all structural constraints while matching the specified attributes in prompts (i.e., num_edges and depth). Figure 12 offers a closer look at 3(a) and 3(b), where each grid point $(X,Y)$ represents accuracy for prompts with $X$ edges and a maximum depth of $Y$. 12(a), 12(b), and 12(c) correspond to the same settings as (baseline), (recursive), and (recursive + instruction) from § 4.1, respectively. The results in Figure 12 show that the recursive generation and instruction tuning improves accuracy across most combinations of (# Edges, Depth). However, some configurations in 12(b) and 12(c) exhibit lower accuracy, likely due to the distribution of training data in terms of edge count and depth.
D.2 Structured Reasoning Results varying Model Sizes
Figure 14 presents the microservice call graph generation accuracy across different model sizes. Overall, models with a larger number of parameters demonstrate higher accuracy, with this trend being particularly evident in 14(c). Notably, models with more parameters perform better as the depth of prompts increases. For instance, the 13B model achieves a 20 percentage point improvement over the 7B model for inputs with a depth greater than 4 as shown in 14(b).
D.3 Memorization
Figure 15: Proportion (%) of synthetic call graphs found in training data varying temperature parameters.
We assess whether synthetic traces are generated by memorizing training data by measuring the percentage of traces that exactly match the structures and call graph attributes found in the training data. Specifically, for the synthetic traces generated in § 4.1, we compute the proportion that exhibit identical call graph structures and edge attributes as those in the training data.
Figure 15 presents the proportion of memorized synthetic traces. Notably, traces generated using the Baseline method exhibit a relatively high level of memorization, with proportions ranging from 16% to 24%. In contrast, our methods (Recursive and Recursive+Instruction) demonstrate significantly lower memorization, with proportions ranging from 3% to 5%.
These results suggest that our recursive generation method is effective not only in producing more structured traces, as shown in § 4.1, but also in minimizing the memorization of training data. This helps generate more diverse synthetic traces.
D.4 More Experiments on Similarity Between Real and Synthetic Traces
@@
D.5 More Experiments on Using Synthetic Traces in ML Use Cases